# ANTS: Shaping the Adaptive Negative Textual Space by MLLM for OOD Detection

### v1

@@ 4.3 Visually Similar Negative Labels
$\mathcal{Y}^{-}_{vsl}=\mathcal{G}_{vsnl}(\mathcal{Y},\mathcal{X}_{neg},f_{mllm},M),$ (9)
@@
$\mathcal{Y}^{\prime}$ $=\text{Top}\left(\mathcal{Y},F(\mathcal{Y}),\delta\right),$ (10)
$\quad F(y_{i})$ $=\frac{|\{{\bm{x}}\in\mathcal{X}_{neg}\mid H({\bm{x}})=y_{i}\}|}{|\mathcal{X}_{neg}|},\quad\forall y_{i}\in\mathcal{Y},$ (11)
where $H({\bm{x}})$ is the CLIP-based ID classifier with ID text features as weight, $|\cdot|$ measures the set size, $F(y_{i})$ represents the proportion of negative images in $\mathcal{X}_{neg}$ being classified as $y_{i}\in\mathcal{Y}$, $F(\mathcal{Y})$ is the collection of $F(y_{i})$, and $\delta\in(0,1)$ serves as the selection ratio.
@@
$\mathcal{Y}^{-}_{vsnl}=\mathcal{G}_{vsnl}(\mathcal{Y}^{\prime},\mathcal{X}_{neg},f_{mllm},M).$ (12)
@@
Algorithm 1 Shaping Adaptive Negative Textual Space by MLLM
2:for each incoming batch $\mathcal{X}_{t}$ do
3:  Filter historical test samples with high $\mathcal{S}_{nl}$ using Eq. 5;
6:  Identify ID label subset similar to negative images using Eq. 9;
9:  Calculate the dynamic weighting $\lambda$ using Eq. 15;
10:  Get final score $S_{ada}$ by weighting the two scores using Eq. 14. //Adaptive Weighted Score
11:end for
12:Output Collected score $S_{ada}$.
@@ 4.4 Adaptive Weighted Score
While expressive negative sentences enhance far-OOD detection by capturing detailed characteristics of negative images, and visually similar negative labels improve near-OOD detection by focusing on OOD-related ID classes, both of them rely on the assumption that the testing scenario (near-OOD or far-OOD) is known beforehand. In real-world applications, this assumption often fails due to the dynamic nature of open environments. To address this, we propose an adaptive weighting strategy to balance these two scoring functions with an adaptive weight $\lambda\in[0,1]$:
@@
where $F(a,b)=\frac{1-a}{(1-a)+(1-b)}\in(0,1)$. One can see that when $\frac{1}{|\mathcal{X}_{neg}|}\sum_{{\bm{v}}\in\mathcal{X}_{neg}^{h}}S_{ens}({\bm{v}})>\frac{1}{|\mathcal{X}_{neg}|}\sum_{{\bm{v}}\in\mathcal{X}_{neg}}S_{vsnl}({\bm{v}})$, $\lambda$ approaches 0; otherwise, $\lambda$ approaches 1. The algorithm is summarized in Alg. 1.
@@
MSP [22] 87.44 58.36 79.73 73.72 79.67 74.41 79.69 71.93 81.63 69.61
ZOC [18] 86.09 87.30 81.20 81.51 83.39 73.06 76.46 98.90 81.79 85.19
CLIPN [72] 95.27 23.94 93.93 26.17 92.28 33.45 90.93 40.83 93.10 31.10
LSN [53] 95.83 21.56 94.35 26.32 91.25 34.48 90.42 38.54 92.26 30.22
LoCoOp [51] 93.93 29.45 90.32 41.13 90.54 44.15 93.24 33.06 92.01 36.95
ID-Like [4] 98.19 8.98 91.64 42.03 90.57 44.00 94.32 25.27 93.68 30.07
TagOOD [40] 98.97 5.00 92.22 29.70 87.81 40.40 90.60 36.31 92.40 27.85
NegPrompt [42] 90.49 37.79 92.25 32.11 91.16 35.52 88.38 43.93 90.57 37.34
SCT [78] 95.86 13.94 95.33 20.55 92.24 29.86 89.06 41.51 93.27 26.47
CoVer [82] 95.98 22.55 93.42 32.85 90.27 40.71 90.14 43.39 92.45 34.88
LAPT [85] 99.63 1.16 96.01 19.12 92.01 33.01 91.06 40.32 94.68 23.40
CMA [33] 99.62 1.65 96.36 16.84 93.11 27.65 91.64 33.58 95.13 19.93
@@
MCM [48] 94.59 32.20 92.25 38.80 90.31 46.20 86.12 58.50 90.82 43.93
CoVer [82] 95.98 22.55 93.42 32.85 90.27 40.71 90.14 43.39 92.45 34.88
EOE [6] 97.52 12.29 95.73 20.40 92.95 30.16 85.64 57.63 92.96 30.09
NegLabel [32] 99.49 1.91 95.49 20.53 91.64 35.59 90.22 43.56 94.21 25.40
OODD [77] 99.36 2.22 95.01 21.49 87.10 44.76 93.27 30.69 93.69 24.79
CLIPScope [20] 99.61 1.29 96.77 15.56 93.54 28.45 91.41 38.37 95.30 20.88
AdaNeg [84] 99.71 0.59 97.44 9.50 94.55 34.34 94.93 31.27 96.66 18.92
CSP [8] 99.60 1.54 96.66 13.66 92.90 29.32 93.86 25.52 95.76 17.51
ANTS 99.69 0.76 98.53 6.50 95.10 27.45 96.38 18.52 97.40 13.31
@@ 5.1 Experiment Setup
Datasets and benchmarks. Following [27], we select ImageNet-1K [12] as the ID dataset and use iNaturalist [68], SUN [75], Places [86], and Textures [10] as the OOD test datasets. We also validate our method on the OpenOOD benchmark, which contains SSB-hard [69] and NINCO [5] as near-OOD datasets, and iNaturalist [68], Texture [10], and OpenImage-O [71] as far-OOD datasets.
Implementation Details. We use the visual encoder of ViT-B/16 pretrained by CLIP [57]. Following NegLabel [32], we adopt the text prompt of ‘The nice <label>.’, set temperature $\tau$ =0.01, and define the number $M$ of negative labels as 5000 with group size 100. Additionally, we set the threshold $\gamma=0.9$ and set $\eta$ = 0.5.
@@ 5.2 Main Results
MCM [48] 79.02 68.54 60.11 84.77
NegLabel [32] 68.18 27.34 76.92 93.30
EOE [6] 82.93 46.73 66.94 89.14
AdaNeg [83] 67.51 17.31 76.70 96.43
@@
We compare our proposed ANTs with other state-of-the-art CLIP-based methods, which can be categorized into two main types: fine-tuning CLIP followed by inference [53, 42, 4, 51, 78, 85, 33], and directly using CLIP for zero-shot inference [18, 55, 32, 6, 8, 77]. To validate that ANTS can better represent the actual OOD space compared to those NL-based methods, we compare ANTS with NegLabel [32], which retrieves negative labels from corpus databases, and EOE [6], which directly generates fixed negative labels through LLMs. In addition, we compare ANTS with those methods utilizing OOD images to obtain image proxies at test time, including OODD [77] and AdaNeg [83].
ImageNet-1k Benchmark. The results are shown in Tab. 1. We can see that other CLIP-based training-based methods [72, 53, 51, 4, 42, 78, 85] learn a negative branch or prompts by synthesizing negative samples, but they often fail to reflect the true OOD space. Our ANTS shapes a more accurate negative space from actual OOD images using MLLMs, and it achieves state-of-the-art performance on the ImageNet-1k benchmark without additional training, emphasizing its exceptional performance in zero-shot scenarios. Other NL-based methods such as NegLabel [32], EOE [6] and CSP [8] directly retrieve negative labels from corpus datasets or generate them using LLMs, but they lack supervision from auxiliary OOD images. ANTS consistently achieves remarkable improvements, which demonstrates the advantages of utilizing the understanding capabilities of MLLMs to shape a more accurate NL space. Compared with other test-time adaptation methods such as AdaNeg [84] and OODD [77], which store test images in memory to calculate the image proxy score and then combine the scores from both modalities, ANTS uses a text-only score to eliminate the modality gap when calculating the OOD score with ID classes, leading to better OOD detection results.
OpenOOD Benchmark. The results are shown in Tab. 2. Though the NL-based methods such as NegLabel [32] and EOE [6] can handle small-scale near-OOD scenarios (e.g., using ImageNet-10 and ImageNet-20 as ID and OOD data, respectively, and the results are presented in the supplementary material), these methods struggle to handle large-scale near-OOD scenarios such as using ImageNet-1k as ID. NegLabel [32] can select negative labels with sufficient semantic differences from ID labels, but it cannot effectively represent near-OOD semantics. EOE [6] uses LLM to generate visually similar labels for all ID classes, bringing a large amount of false negative labels with the increase in the number of ID classes. Our proposed ANTS first identifies a subset of ID classes similar to OOD images, as shown in Fig. 6(a), then leverages the reasoning capabilities of MLLMs to generate visually similar labels. As a result, ANTS significantly outperforms its closest competitors [32, 6, 84] in both near-OOD and far-OOD scenarios, validating its scalability.
@@ 5.3 Analyses and Discussions
(a) ID noise rate of negative images
(b) Selection ratio $\delta$
(c) Weight $\lambda$
(d) Backbones
(e) MLLMs prompts
(f) MLLMs
Figure 7: Analysis on (a) length of negative text, (b) selection ratio $\delta$, (c) weight $\lambda$, (d) CLIP image encoder backbones, (e) MLLMs prompts, and (f) different MLLMs. We use Texture [10] and NINCO [5] datasets as Far-OOD and Near-OOD, respectively.
@@
Robustness of ANTS to Noisy Negative Images. We investigate the impact of different rates of ID noise in negative images on the model OOD detection performance, as shown in Fig. 7(a). When the ID ratio of negative images is relatively low (less than 0.5), our ANTS achieves excellent OOD detection performance across different ID ratios. This demonstrates the robustness of ANTS against ID noise.
Ratio $\delta$. As shown in Fig. 7(b), generating visually similar labels for all ID classes ( $\delta=1$) performs poorly due to numerous false negative labels, as illustrated in Fig. 6(a). However, using a too small $\delta$ will fail to adequately cover the OOD distribution. We set $\delta=0.08$ in all experiments, although it is not optimal for specific datasets.
Weight $\lambda$. As shown in Fig. 7(c), a larger $\lambda$ emphasizes the $S_{ens}(x)$ score, improving far-OOD detection, while a smaller $\lambda$ prioritizes the $S_{vsnl}(x)$ score, enhancing near-OOD detection. Our adaptive strategy automatically selects a suitable $\lambda$ for various OOD settings.
Different Backbones. As illustrated in Fig. 7(d), larger visual backbones generally achieve improved OOD detection. Besides, our ANTS can generalize well to various VLM backbones, demonstrating its robustness.
Different Prompts for MLLMs. To investigate the effectiveness of the prompts for ENS and VSNL, we designed two additional prompts: an inexpressive prompt and a dissimilar prompt. Specifically, the inexpressive prompt asks MLLMs to generate descriptions for negative images using less than three words. The dissimilar prompt prompts MLLMs to generate negative categories that are visually dissimilar to the high frequency ID classes. As shown in Fig. 7(e), OOD detection performance degrades with these two instructions, demonstrating the effectiveness of our proposed prompts.
Various MLLMs. To investigate the impact of different MLLMs in adaptive negative textual space construction, we conducted experiments using MLLMs with varying parameter sizes, as illustrated in Fig. 7(f). Models of different sizes show little variation in performance on far OOD. However, in the near-OOD settings, larger MLLMs perform better due to their stronger reasoning capabilities and LLaVA-7B achieves the best performance.

### v3

@@ 4.3 Visually Similar Negative Labels
$\mathcal{Y}^{-}_{vsl}=\mathcal{G}_{vsnl}(\mathcal{Y},\mathcal{X}^{\rm his}_{test},f_{mllm},M),$ (9)
@@
$F(y_{i})$ $=\frac{|\{{\bm{x}}\in\mathcal{X}^{\rm his}_{test}\mid H({\bm{x}})=y_{i}\}|}{|\mathcal{X}^{\rm his}_{test}|},\quad\forall y_{i}\in\mathcal{Y},$ (10)
$\quad\mathcal{Y}^{\prime}$ $=\text{Top}\left(\mathcal{Y},F(\mathcal{Y}),\delta\right),$ (11)
where $H({\bm{x}})$ is the CLIP-based ID classifier with ID text features as weight, $|\cdot|$ measures the set size, $F(y_{i})$ represents the proportion of historical test images in $\mathcal{X}^{\rm his}_{test}$ being classified as $y_{i}\in\mathcal{Y}$, $F(\mathcal{Y})$ is the collection of $F(y_{i})$, and $\delta\in(0,1)$ serves as the selection ratio.
@@
$\mathcal{Y}^{-}_{vsnl}=\mathcal{G}_{vsnl}(\mathcal{Y}^{\prime},\mathcal{X}^{\rm his}_{test},f_{mllm},M).$ (12)
@@ 4.4 Adaptive Weighted Score
Existing OOD detection methods rely on the assumption that the testing scenario (near-OOD or far-OOD) is human defined beforehand, but in real-world applications, this assumption often fails due to the dynamic nature of open environments. To address this, we propose an adaptive weighting strategy to balance these two scoring functions with an adaptive weight $\lambda\in[0,1]$:
@@
where $F(a,b)=\frac{1-a}{(1-a)+(1-b)}\in(0,1)$. One can see that when $\frac{1}{|\mathcal{X}_{neg}|}\sum_{{\bm{v}}\in\mathcal{X}_{neg}}S_{ens}({\bm{v}})>\frac{1}{|\mathcal{X}_{neg}|}\sum_{{\bm{v}}\in\mathcal{X}_{neg}}S_{vsnl}({\bm{v}})$, $\lambda$ approaches 0; otherwise, $\lambda$ approaches 1.
@@
MSP [hendrycks2016baseline] 87.44 58.36 79.73 73.72 79.67 74.41 79.69 71.93 81.63 69.61
ZOC [esmaeilpour2022zero] 86.09 87.30 81.20 81.51 83.39 73.06 76.46 98.90 81.79 85.19
CLIPN [wang2023clipn] 95.27 23.94 93.93 26.17 92.28 33.45 90.93 40.83 93.10 31.10
LSN [nie2024out] 95.83 21.56 94.35 26.32 91.25 34.48 90.42 38.54 92.26 30.22
LoCoOp [miyai2024locoop] 93.93 29.45 90.32 41.13 90.54 44.15 93.24 33.06 92.01 36.95
ID-Like [bai2024id] 98.19 8.98 91.64 42.03 90.57 44.00 94.32 25.27 93.68 30.07
NegPrompt [li2024learning] 90.49 37.79 92.25 32.11 91.16 35.52 88.38 43.93 90.57 37.34
SCT [yu2024self] 95.86 13.94 95.33 20.55 92.24 29.86 89.06 41.51 93.27 26.47
LAPT [zhang2024lapt] 99.63 1.16 96.01 19.12 92.01 33.01 91.06 40.32 94.68 23.40
CMA [kim2025enhanced] 99.62 1.65 96.36 16.84 93.11 27.65 91.64 33.58 95.13 19.93
@@
MCM [ming2022delving] 94.59 32.20 92.25 38.80 90.31 46.20 86.12 58.50 90.82 43.93
CoVer [zhang2024if] 95.98 22.55 93.42 32.85 90.27 40.71 90.14 43.39 92.45 34.88
EOE [cao2024envisioning] 97.52 12.29 95.73 20.40 92.95 30.16 85.64 57.63 92.96 30.09
NegLabel [jiang2024negative] 99.49 1.91 95.49 20.53 91.64 35.59 90.22 43.56 94.21 25.40
OODD [yang2025oodd] 99.36 2.22 95.01 21.49 87.10 44.76 93.27 30.69 93.69 24.79
AdaNeg [zhang2024adaneg] 99.71 0.59 97.44 9.50 94.55 34.34 94.93 31.27 96.66 18.92
CSP [chen2024conjugated] 99.60 1.54 96.66 13.66 92.90 29.32 93.86 25.52 95.76 17.51
SynOOD [li2025synthesizing] 99.57 1.57 95.82 20.46 97.37 12.12 95.29 22.94 97.01 14.27
ANTS 99.75 0.54 98.77 5.43 96.10 20.21 96.38 18.52 97.75 11.20
Algorithm 1 Shaping Adaptive Negative Textual Space by MLLM
2:for each incoming batch $\mathcal{X}_{t}$ do
3:  Filter historical test samples with $\mathcal{S}_{nl}$ using Eq. 5;
6:  Identify ID label subset similar to historical test images using Eq. 9;
9:  Calculate the dynamic weighting $\lambda$ using Eq. 15;
10:  Get the final score $S_{ada}$ by weighting the two scores using Eq. 14. //Adaptive Weighted Score
11:end for
12:Output Collected score $S_{ada}$.
@@ 5.1 Experiment Setup
Datasets and benchmarks. Following [huang2021importance], we select ImageNet-1K [deng2009imagenet] as the ID dataset and use iNaturalist [van2018inaturalist], SUN [xiao2010sun], Places [zhou2017places], and Textures [cimpoi2014describing] as the OOD test datasets. We also validate our method on the OpenOOD benchmark, which contains SSB-hard [vaze2021open] and NINCO [bitterwolf2023or] as near-OOD datasets, and iNaturalist [van2018inaturalist], Texture [cimpoi2014describing], and OpenImage-O [wang2022vim] as far-OOD datasets.
Implementation Details. We use the visual encoder of ViT-B/16 pretrained by CLIP [radford2021learning]. We adopt the LLaVA-1.5-7B model as the default MLLM for our research, Following NegLabel [jiang2024negative], we adopt the text prompt of ‘The nice <label>.’, set temperature $\tau$ =0.01, and define the number $M$ of negative labels as 10000 with group size 100. Additionally, we set the initial threshold $\gamma=0.9$ and set $\eta$ = 0.5. Our method follows a test-time adaptation setting as [zhang2024adaneg].
@@ 5.2 Main Results
MCM [ming2022delving] 79.02 68.54 60.11 84.77
NegLabel [jiang2024negative] 68.18 27.34 76.92 93.30
EOE [cao2024envisioning] 82.93 46.73 66.94 89.14
AdaNeg [zhang2022out] 67.51 17.31 76.70 96.43
SynOOD [li2025synthesizing] 71.68 17.11 77.55 96.21
@@
Table 3: OOD detection performance on other ID datasets.
[…]
CUB-200-2011 NegLabel [jiang2024negative] 99.93 0.13
ANTS (Ours) 99.95 0.01
STANFORD-CARS NegLabel [jiang2024negative] 99.99 0.01
ANTS (Ours) 99.99 0.00
Food-101 NegLabel [jiang2024negative] 99.90 0.40
ANTS (Ours) 99.92 0.05
Oxford-IIIT Pet Neglabel [jiang2024negative] 99.62 1.70
ANTS (Ours) 99.99 0.02
ImageNet-1k Benchmark. The results are shown in Tab. 1. We can see that other CLIP-based training-based methods [wang2023clipn, nie2024out, miyai2024locoop, bai2024id, li2024learning, yu2024self, zhang2024lapt] learn a negative branch or prompts by synthesizing negative samples, but they often fail to reflect the true OOD space. Other NL-based methods [jiang2024negative, cao2024envisioning, chen2024conjugated] directly retrieve negative labels from corpus datasets or generate them using LLMs, but they lack supervision from auxiliary OOD images. ANTS consistently achieves remarkable improvements, which demonstrates the advantages of utilizing the understanding capabilities of MLLMs to shape a more accurate NL space. Compared with other test-time adaptation methods  [zhang2024adaneg, yang2025oodd], which store test images in memory to calculate the image proxy score and then combine the scores from both modalities, ANTS uses a text-only score to eliminate the modality gap when calculating the OOD score with ID classes, leading to better OOD detection results.
OpenOOD Benchmark. The results are shown in Tab. 2. Though the NL-based methods [jiang2024negative, cao2024envisioning] can handle small-scale near-OOD scenarios (e.g., using ImageNet-10 and ImageNet-20 as ID and OOD data, respectively, and the results are presented in the supplementary material), these methods that selects semantically distant negative labels struggle to handle large-scale near-OOD scenarios such as using ImageNet-1k as ID. However, EOE [cao2024envisioning], while using LLMs to generate visually similar labels, suffers from a growing number of false negatives with increasing ID classes. ANTS first identifies a subset of ID classes similar to OOD images, as shown in Fig. 6(a), then it leverages the reasoning capabilities of MLLMs to generate visually similar labels. As a result, ANTS significantly outperforms its closest competitors [jiang2024negative, cao2024envisioning, zhang2024adaneg] in both near-OOD and far-OOD scenarios, validating its scalability.
Results of other ID datasets. As shown in Tab. 3, our ANTS consistently surpasses existing methods in zero-shot OOD detection method NegLabel[jiang2024negative] across all in-distribution (ID) datasets. We also validate the robustness of ANTS to Domain Shift and Adversarial Examples, the detailed results are available in the supplementary materials.
@@ 5.3 Analyses and Discussions
(a) Initial OOD detector
(b) The lengths of negative sentences.
(c) Selection ratio $\delta$
(d) Weight $\lambda$
(e) Backbones
(f) MLLMs prompts
(g) MLLMs
(h) Temporal shift
Figure 7: Analysis on (a) different initial OOD detectors, (b) the lengths of negative sentences, (c) selection ratio $\delta$, (d) weight $\lambda$, (e) CLIP image encoder backbones, (f) MLLMs prompts, and (g) different MLLMs. (h) Temporal shift. We use Texture [cimpoi2014describing] and NINCO [bitterwolf2023or] datasets as Far-OOD and Near-OOD, respectively.
@@
Analyses on initial OOD detectors. Besides NegLabel, we also tested two other variants: (1) a weak MCM detector, and (2) a cosine-distance filter that selects negative far from ID labels in the feature space. As shown in Tab.  7(a), even with these weaker detectors, our method still outperforms previous SOTA baseline.
Robustness to noisy negative images. The results are presented in the supplementary material.
Analyses on the lengths of negative sentences. Due to the hallucination issue in MLLMs, we analyzed the length of generated negative sentences. As shown in Fig. 7(b), appropriate increases in length enhance expressiveness and OOD detection, while excessive text introduces less discriminative words and hinders performance. Our ENS achieves an optimal balance at an average length of 8.4.
Ratio $\delta$. As shown in Fig. 7(c), generating visually similar labels for all ID classes ( $\delta=1$) performs poorly due to numerous false negative labels, as illustrated in Fig. 6(a). However, using a too small $\delta$ will fail to adequately cover the OOD distribution. We set $\delta=0.08$ in all experiments, although it is not optimal for specific datasets.
Weight $\lambda$. As shown in Fig. 7(d), a larger $\lambda$ emphasizes the $S_{ens}(x)$ score, improving far-OOD detection, while a smaller $\lambda$ prioritizes the $S_{vsnl}(x)$ score, enhancing near-OOD detection. Our adaptive strategy automatically selects a suitable $\lambda$ for various OOD settings.
Different Backbones. As illustrated in Fig. 7(e), larger visual backbones generally achieve improved OOD detection. Besides, our ANTS can generalize well to various VLM backbones, demonstrating its robustness.
Different Prompts for MLLMs. We designed two alternative prompts: an inexpressive prompt that limits negative image descriptions to under three words, and a dissimilar prompt that requests negative categories visually distinct from high-frequency ID classes. As shown in Fig. 7(f), OOD detection performance declines with both alternatives, confirming our proposed prompts’ efficacy.
Various MLLMs. When constructing the adaptive negative space, MLLMs of all sizes showed comparable far-OOD detection, but larger models excelled in near-OOD settings. As shown in Fig. 7(g), LLaVA-7B performed best, owing to stronger reasoning.
Analysis of Temporal Shift. We evaluated ANTS under different temporal shifts. As shown in Fig. 7(h), "Near-to-Far" and "Far-to-Near" indicate testing first on near (or far) OOD, then on the opposite, ANTS maintains strong performance, demonstrating robustness to temporal shift.